In the course of doing valuable biomedical research one day, let's say you make a salt solution by adding 1 mol of NaCl to 90 g of water. 


What is the percent by mass (i.e. mass %) of the solution? 


1
Expert's answer
2022-03-31T08:59:02-0400

Mass of 1 mol of NaCl:

"m=\nu\cdot M=1mol\cdot58.5g/mol=58.5g - NaCl"

Total mass: "m_0=58.5g+90g=148.5g"

Mass % of NaCl: "58.5g/148.5g\cdot 100" %"=39.39" %

Mass % of water: "(100-39.39)" %"=60.61" %